# Correctness Assessment of Code Generated by Large Language Models Using Internal Representations

### v1

@@ 6 Related Work
Hallucinations, where LLMs produce plausible yet incorrect or non-functional code, are another prominent issue in practical code generation [67]. Understanding and addressing these hallucinations is essential for ensuring the reliability of LLMs in software development. Furthermore, research has shown that AI-generated code often contains vulnerabilities, necessitating a deeper investigation into its safety implications [68]. For instance, comparative studies have assessed the security vulnerabilities of ChatGPT-generated code against traditional sources like StackOverflow, revealing notable differences in the safety and correctness of the generated code [69]. Efforts have also been made to analyze the cognitive processes of LLMs during code generation. A study explored whether LLMs pay attention to code structures in a manner similar to human programmers, providing insights into their decision-making and error tendencies [70]. Additionally, uncertainty analysis for LLMs has been proposed to measure their confidence in generated outputs, which could aid in identifying potentially unreliable code [71].
@@
More recent approaches integrate pre-trained models like CodeBERT [40] and CodeT5 [41], which have demonstrated significant effectiveness in both function- and statement-level vulnerability detection tasks. For example, LineVul[90] and LineVD[91] leverage CodeBERT to encode code representations, outperforming traditional methods like IVDetect in identifying vulnerabilities at finer levels of granularity. The success of these pre-trained models is largely attributed to their ability to capture rich contextual information and encode both syntactic and semantic characteristics of code efficiently. As a result, models like CodeBERT and CodeT5 have established themselves as powerful baselines for advancing the field of bug and vulnerability detection, driving progress toward more accurate and effective software quality assurance tools.
@@ 7 Conclusion
In this work, we introduced Openia, a novel framework that leverages the internal representations of code LLMs to assess the correctness of generated code. Unlike traditional black-box approaches, Openia adopts a white-box methodology, systematically analyzing the intermediate signals of code-specific LLMs such as DeepSeek-Coder, CodeLlama, and MagicCoder. Our empirical findings highlight that these internal signals capture latent information, including key programming principles, which strongly correlate with the reliability of the generated code. Openia not only provides a more nuanced and reliable evaluation of code correctness but also demonstrates adaptability and robustness across varying benchmarks and tasks. By unlocking the potential of in-process signals, Openia bridges the gap between code generation and quality assurance, offering a proactive approach to enhancing the reliability of LLM-generated code. Our framework sets the foundation for future research into leveraging the latent capabilities of LLMs to improve AI-driven software development.

### v2

@@ 6 Related Work
Hallucinations, where LLMs produce plausible yet incorrect or non-functional code, are another prominent issue in code generation [70]. Understanding and addressing these hallucinations is essential for ensuring the reliability of LLMs in software development. Furthermore, research has shown that AI-generated code often contains vulnerabilities, necessitating a deeper investigation into its safety implications [71]. For instance, comparative studies have assessed the security vulnerabilities of ChatGPT-generated code against traditional sources like StackOverflow, revealing notable differences in the safety and correctness of the generated code [72]. Efforts have also been made to analyze the cognitive processes of LLMs during code generation. A study explored whether LLMs pay attention to code structures in a manner similar to human programmers, providing insights into their decision-making and error tendencies [73]. Uncertainty analysis for LLMs has been proposed to measure their confidence in generated outputs, which could aid in identifying potentially unreliable code [74].
@@
More recent approaches integrate pre-trained models like CodeBERT [44] and CodeT5 [45], which have demonstrated significant effectiveness in both function-level and statement-level vulnerability detection tasks. For example, LineVul [93] and LineVD [94] leverage CodeBERT to encode code representations, outperforming traditional methods like IVDetect in identifying vulnerabilities at finer levels of granularity. The success of these pre-trained models is largely attributed to their ability to capture rich contextual information and encode both syntactic and semantic characteristics of code efficiently. As a result, models like CodeBERT and CodeT5 have established themselves as powerful baselines for advancing the field of bug and vulnerability detection, driving progress toward more accurate and effective software quality assurance tools.
@@
While our work and prior studies [37, 39, 97, 99] similarly leverage internal representations of LLMs to detect hallucination, Openia fundamentally differs in both task objective and methodological approach. These exiting studies [37, 39, 97, 99] primarily focus on hallucination detection in NLG, where outputs are typically unstructured and correctness is subjective, relying on human judgment or alignment with external factual knowledge. In contrast, Openia targets code generation, in which correctness is objectively defined through syntax, logic, and runtime behavior, as can be verified by functional test cases. Moreover, due to the nature of NLG tasks, prior works mainly investigate semantic consistency or factual alignment within model responses. Azaria et al. [97] detect hallucinations by training a classifier on the internal representations of both true and false factual statements. INSIDE  [39] measures the semantic diversity and consistency across multiple responses to the same question to assess the risk of hallucination. However, these techniques are not directly applicable to code. Unlike natural language, where factual inconsistency often indicates hallucination, a programming task may have multiple correct solutions that differ in implementation details or algorithms. As a result, inconsistency across generated responses does not necessarily imply incorrectness. This key difference highlights the need for a specific approach for code generation assessment like Openia.
@@
In general, these white-box approaches [38, 39, 97, 37], leverage internal neural signals of LLMs such as hidden layers, attention weights, or token embeddings, etc. to explore the reasoning process of LLMs during inference. This reflects a model-centric perspective, which seeks to understand how the model internally “thinks” and makes decisions, even when its reasoning is not explicitly expressed in the output. Meanwhile, chain-of-thought (CoT) prompting is an output-centric technique that encourage the model to produce step-by-step reasoning as part of its response [103]. CoT treats this verbalized reasoning as a proxy for the model’s internal thought process. While both approaches aim to uncover how LLMs arrive at their outputs, they differ in whether reasoning is inferred from internal representations or observed directly in the generated response.
@@ 7 Conclusion
This work introduces Openia, a novel framework that leverages the internal representations of code LLMs to assess the correctness of generated code. Unlike traditional black-box approaches, Openia adopts a white-box methodology, systematically analyzing the intermediate states of specialized code LLMs such as DeepSeek Coder, Code Llama, and Magicoder. Our empirical findings demonstrate that these internal signals capture latent information, including fundamental programming principles, which strongly correlate with the reliability of the generated code. Openia not only provides a more nuanced and reliable evaluation of code correctness but also exhibits strong adaptability and robustness across diverse benchmarks and tasks. By unlocking the potential of in-process signals, Openia bridges the gap between code generation and quality assurance, offering a proactive approach to enhancing the reliability of LLM-generated code. Our framework sets the foundation for future research into leveraging the latent capabilities of LLMs to improve AI-driven software development.
Acknowledgement
This research is supported by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 102.03-2023.14. This research is also partly supported by OpenAI’s Researcher Access Program.